An object is thrown vertically and has an upward velocity of 25 m/s


when it reaches one fourth of its maximum height above its launch point.


What is the initial (launch) speed of the object?



Please explain how to answer this exercise.



1
Expert's answer
2021-12-16T11:31:34-0500

The velocity of an object thrown upward decreases with height according to the following law:


"h=\frac{H_\text{max}}4=\frac{v_f^2-v_i^2}{2g},\\\space\\
v_i^2=v_f^2-2gh=25^2-2·(-9.8)·\frac{H_\text{max}}4."

The maximum height:


"H_\text{max}=\frac{v_f^2-v_i^2}{2g}=\frac{0-v_i^2}{2g}=\frac{-v_i^2}{2g}."

Substitute this into the equation above:


"v_i^2=25^2-2·(-9.8)·\frac{-v_i^2}{4·2·(-9.8)},\\\space\\
v_i^2=\frac{25^2}{1-1/4}=833\text{ (m/s})^2,\\\space\\
v_i=28.9\text{ m/s}."